A stock had returns of 8%, -13%, -7%, 29% for four of the last 5 years. If the average return of the stock over this period was 11%, what was the stock’s return for the missing year?
Answer #1

Average return=Total return/Total time period

11=(8-13-7+29+return for the missing year)/5

(11*5)=17+return for the missing year

return for the missing year=55-17

=38%